# Observational properties of black hole in quantum fluctuation modified gravity

### v1

@@ 1 Introduction
Various observations have confirmed that the expansion of the universe is accelerating. Modified gravity theories, such as $f(R)$ gravity, $f(R,T)$ theory, and Gauss-Bonnet gravity, are popular attempts to explain this phenomenon. If considering Heisenberg’s nonperturbative quantization [1, 2, 3]: the metric is decomposed into the sum of the classical part and the quantum fluctuation part, the corresponding quantum Einstein gravity engenders modified gravity model at the classical level, which can be called as quantum fluctuation modified gravity (QFMG). QFMG has been realized in some special models, see for example [4, 5, 6, 7, 8]. QFMG has some interesting properties, see for example, it will result in a non-minimal coupling between geometry and matter, which gives an possible microscopic quantum description of the matter creation processes in $f(R,T)$ or $f(R,L_{\rm{m}})$ gravity [5, 9]. Baryogenesis in QFMG was recently discussed in [10]. Black hole (BH) solutions in QFMG were considered in [9, 11]. Friedmann-Lemaitre-Robertson-Walker cosmology in QFMG was investigated in [4, 5, 6, 12, 7].
@@
The observations of the shadows of the supermassive BHs at the centers of the galaxy M87* and the Milky Way by the Event Horizon Telescope (EHT) Collaboration [13, 14], has given a boost to the study of BH physics. BHs are usually surrounded by thin accretion disk. Due to the strong gravitational field, the thin accretion disk will radiate high-energy fluxes [15, 16, 17, 18, 19, 20, 21, 22], which have significant effects on BH’s shadow. The study of BH’s shadow has been ongoing for a long time. In 1979, Luminet had obtained simulated images of a BH with thin accretion disk, illustrating that the BH’s shadow depends on the accretion flow [23]. The shadow for Kerr BH with Keplerian accretion disk had been discussed in detail in [24]. Using the general relativistic ray-tracing code, reference [25] had investigated the shadow images of a geometrically thick and infinite thin accretion disk around the compact objects. By considering the shape and the size of BH’s shadow depending on the spacetime metric, the observation of BH provides an important way to extract useful information about the strong gravitational field, and helps us to distinguish various gravitational theories and even to investigate the quantum gravity effects. Recently a large nunber of studies on BH shadow have been made in various BH spacetimes, see for example [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55].
@@
BH are important research subjects in gravitational theories, serving as a crucial test for our understanding of general relativity (GR) and modified gravity. Because of the presence of the matter-geometry coupling in QFMG, it is expected differences between BH solutions in GR and QFMG. Such differences may become more prominent for strong gravitational field. So it is important to investigate the effects imposed by QFMG in the scale of compact objects, such as BHs. In [11], an interesting BH solution of QFMG has been obtained, which has two characteristic parameters besides mass: one parameter characterizes the quantum fluctuation of metric, the other characterizes the matter surrounded the BH. Here we will discuss the effects of these two parameters on the thin accretion disk and the shadow of BH.
@@
The paper is organized as follows. In Section II, we will briefly review the line element and model the motion of massive and massless particles respectively. In Section III we model thin accretion disk in the geometry of the QFMGBH. In section IV, we use GYOTO, a general relativistic ray-tracing code [11] to simulate the image of the QFMGBH’s shadow. Finally in section VI we summarize the results and their implications. Throughout the article we use geometrized units setting $c=1$ with the metric signature ,+,+,+-,+,+,+- , + , + , +.
2 The spacetime and the motion of particles
@@
2.1 Black hole in quantum fluctuation modified gravity
In [11], an interesting QFMGBH solution for the gravitational field equations in QFMG proposed in [4] was obtained, which reduces to different classes of BHs surrounded by Kiselev fluids, by taking some specific values of the parameter of the equation of state (EoS). The geometry of the static and spherically symmetric QFMGBH in Schwarzschild coordinates is given by
$\begin{split}ds^{2}&=-g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}\\ &=-Bdt^{2}+\frac{1}{B}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2},\end{split}$ (2.1)
@@ 2.1 Black hole in quantum fluctuation modified gravity
$\begin{split}B&=1-\frac{2M}{r}+Dr^{-\frac{2(1+3\omega-4\omega\alpha)}{2-3\alpha+\omega\alpha}},\\ \end{split}$ (2.2)
where $D$ is an integration constant, the parameter $\alpha$ characterizing the fluctuation of metric, and the parameter $\omega$ denotes the EoS of Kiselev fluid. If there are no fluctuations and fluids, solution (2.2) will reduce to the Schwarzschild BH. Here, we mainly focus on the impacts of parameters $\omega$ and $\omega$ on the research objects. For the convenience of calculations, we take $D=\lambda M$. The dimension of $\lambda$ depends on the parameters $\omega$ and $\alpha$. When the BH is surrounded by different fluids, such as radiation, dust, quintessence, cosmological constant, or phantom, solution (2.2) reduces to different BH spacetime.
2.2 Asymptotically flat condition
@@
First we assume that the QFMGBH spacetime should be asymptotically flat at infinity, namely, it has to satisfy this condition $B\to 1$ when $r\to\infty$. In [4], the range of the parameter $\alpha$ is $(-1,1)$. We can further constrain $\alpha$ by the asymptotically flat condition, The relevant inequality is shown as
$P=-\frac{2(1+3\omega-4\omega\alpha)}{2-3\alpha+\omega\alpha}<0.$ (2.3)
Through analyzing the inequality (2.3), we can know that as $\alpha$ varies in range $(-1,1)$, the $P$ identically equals to $1$ for $\omega=-1$, which doesn’t satisfy the asymptotically flat condition. So we don’t consider this case ($\omega=-1$) further. In the same way, we can acquire the ranges of $\alpha$ for other different values of $\omega$: the range of $\alpha$ for $\omega=1/3$ is $(-1,3/4)$, for $\omega=0$ is $(-1,2/3)$, for $\omega=-2/3$ is $(3/8,6/11)$, and for $\omega=-4/3$ is $(6/13,9/16)$. We visualize these ranges in figure 1.
@@
2.3 Stable circular orbits
From the definition of the BH, we can know that the BH cannot be directly observed, but we can understand the properties of the central objects by researching the properties of its surroundings [56], so the time-like geodesics around compact astrophysical objects are very crucial. As usually, we assume that the accretion disk is located in the equatorial plane of the BH ($\theta=\pi/2$). The Lagrangian of particles moving in the equatorial plane of a static spherically symmetric BH is as follows
$\begin{split}\mathcal{L}&=\frac{1}{2}g_{\mu\nu}\dot{x^{\mu}}\dot{x^{\nu}}\\ &=\frac{1}{2}g_{tt}\dot{t}^{2}+\frac{1}{2}g_{rr}\dot{r}^{2}+\frac{1}{2}g_{\phi\phi}\dot{\phi}^{2},\end{split}$ (2.4)
@@ 2.3 Stable circular orbits
$\begin{split}g_{tt}\dot{t}&=-E,\\ g_{\phi\phi}\dot{\phi}&=L,\end{split}$ (2.5)
@@
$\begin{split}\dot{t}&=-\frac{E}{g_{tt}},\\ \dot{\phi}&=\frac{L}{g_{\phi\phi}}.\end{split}$ (2.6)
@@
$V_{\rm{eff}}(E,L,r)=-1+\frac{g_{\phi\phi}E^{2}+g_{tt}L^{2}}{-g_{tt}g_{\phi\phi}},$ (2.7)
where $V_{\rm{eff}}=g_{rr}\dot{r}^{2}$ denotes the effective potential of particles. Circular orbits exist at positions where $\dot{r}=0$, which corresponds to the conditions $V_{\rm{eff}}=0$ and $V_{\rm{eff}}^{\prime}=0$ [57, 58, 59], where the prime denotes differentiation with respect to $r$. Under these conditions, we can derive expressions for the specific energy $E$, the specific angular momentum $L$, and the specific angular velocity $\Omega$ of particles in the equatorial plane of a Schwarzschild-like black hole spacetime.
$\Omega=\frac{d\phi}{dt}=\frac{\sqrt{-g_{tt,r}g_{\phi\phi,r}}}{g_{\phi\phi,r}},$ (2.8)
$E=-\frac{g_{tt}}{\sqrt{-g_{tt}-g_{\phi\phi}\Omega^{2}}},$ (2.9)
$L=\frac{g_{\phi\phi}\Omega}{\sqrt{-g_{tt}-g_{\phi\phi}\Omega^{2}}}.$ (2.10)
We can derive the innermost stable circular orbit (ISCO) for particles in the circular orbits by putting the Eqs. (2.8)–(2.10) into $V_{\rm{eff}}^{\prime\prime}=0$.
3 Physical properties of thin accretion disk
@@
Firstly, we provide a brief overview of the key properties of the accretion disk in the Novikov-Thorne model [60, 61]. $(\textbf{1})$ The spacetime around the central mass object is static, axisymmetric, and asymptotically flat. $(\textbf{2})$ The gravitational effect that the accretion disk generate can be ignored. $(\textbf{3})$ The accretion disk is geometrically thin and optically thick, with the disk’s semi-thickness $h$ being significantly smaller than its radial extent $r$. As a result, light is absorbed when it passes vertically through the disk. $(\textbf{4})$ The inner edge of the disk is defined by the radius of the marginally stable orbit, with particles confined between the ISCO and the outer edge. $(\textbf{5})$ The accretion disk coincides with the equatorial plane of the compact accreting object. $(\textbf{6})$ The radiation emitted from the disk is modeled as a black-body spectrum. $(\textbf{7})$ The mass accretion rate $\dot{M}$ is constant over time.
3.1 The radiant energy flux
Considering the accretion disk around a QFMGBH, the radiant energy flux on the surface of the accretion disk is given by [61, 62]
[…]
The integral above is only available in cylindrical coordinates. For spherical coordinates, we need to utilize [63]
[…]
where $\dot{M_{0}}$ is the accretion rate of mass.

### v2

@@ I Introduction
Various observations have confirmed that the expansion of the universe is accelerating. Modified gravity theories, such as $f(R)$ gravity, $f(R,T)$ theory, and Gauss-Bonnet gravity, are popular attempts to explain this phenomenon. If considering Heisenberg’s nonperturbative quantization (Dzhunushaliev:2013nea, ; Dzhunushaliev:2015mva, ; Dzhunushaliev:2015hoa, ): the metric is decomposed into the sum of the classical part and the quantum fluctuation part, the corresponding quantum Einstein gravity engenders modified gravity model at the classical level, which can be called as quantum fluctuation modified gravity (QFMG). QFMG has been realized in some special models, see for example Yang:2015jla ; Liu:2016qfx ; Bernardo:2021ynf ; Chen:2021oal ; Lima:2023qdv . QFMG has some interesting properties, see for example, it will result in a non-minimal coupling between geometry and matter, which gives an possible microscopic quantum description of the matter creation processes in $f(R,T)$ or $f(R,L_{\rm{m}})$ gravity Liu:2016qfx ; Yang:2020lxv . Baryogenesis in QFMG was recently discussed in Yang:2024gnf . Black hole (BH) solutions in QFMG were considered in Yang:2020lxv ; Hua:2024jzx . Friedmann-Lemaitre-Robertson-Walker cosmology in QFMG was investigated in Yang:2015jla ; Liu:2016qfx ; Bernardo:2021ynf ; Haghani:2021iqe ; Chen:2021oal .
@@
The observations of the shadows of the supermassive BHs at the centers of the galaxy M87* and the Milky Way by the Event Horizon Telescope (EHT) Collaboration EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp , has given a boost to the study of BH physics. BHs are usually surrounded by thin accretion disk. Due to the strong gravitational field, the thin accretion disk will radiate high-energy fluxes Collodel:2021gxu ; Liu:2024brf ; Wu:2024sng ; Das:2024tyh ; Abbas:2023rzk ; He:2022lrc ; Gyulchev:2019tvk ; Gyulchev:2020cvo , which have significant effects on BH’s shadow. The study of BH’s shadow has been ongoing for a long time. In 1979, Luminet had obtained simulated images of a BH with thin accretion disk, illustrating that the BH’s shadow depends on the accretion flow Luminet:1979nyg . The shadow for Kerr BH with Keplerian accretion disk had been discussed in detail in Beckwith:2004ae . Using the general relativistic ray-tracing code, reference Vincent_2011 had investigated the shadow images of a geometrically thick and infinite thin accretion disk around the compact objects. By considering the shape and the size of BH’s shadow depending on the spacetime metric, the observation of BH provides an important way to extract useful information about the strong gravitational field, and helps us to distinguish various gravitational theories and even to investigate the quantum gravity effects. Recently a large nunber of studies on BH shadow have been made in various BH spacetimes, see for example Vagnozzi:2022moj ; Uniyal:2022vdu ; Heydari-Fard:2022xhr ; Zare:2024dtf ; Ma:2022jsy ; Meng:2024puu ; Fu:2022yrs ; Mirzaev:2022xpz ; Zhang:2024jrw ; Zhang:2024lsf ; Gyulchev:2024iel ; Li:2022eue ; Zhu:2024vxw ; Luo:2024nul ; Rodriguez:2024ijx ; Macedo:2024qky ; Hu:2023pyd ; Meng:2023uws ; Johnson:2023skw ; Tang:2023lmr ; Fauzi:2024nta ; Zheng:2024ftk ; Yang:2024nin ; Gao:2023mjb ; Zhang:2023okw ; Vincent:2022fwj ; Ban:2024qsa ; Sui:2023tje ; Benavides-Gallego:2024jyw ; Liu:2024lbi .
@@
BH are important research subjects in gravitational theories, serving as a crucial test for our understanding of general relativity (GR) and modified gravity. Because of the presence of the matter-geometry coupling in QFMG, it is expected differences between BH solutions in GR and QFMG. Such differences may become more prominent for strong gravitational field. So it is important to investigate the effects imposed by QFMG in the scale of compact objects, such as BHs. In Hua:2024jzx , an interesting BH solution of QFMG has been obtained, which has two characteristic parameters besides mass: one parameter characterizes the quantum fluctuation of metric, the other characterizes the matter surrounded the BH. Here we will discuss the effects of these two parameters on the thin accretion disk and the shadow of BH.
@@
The paper is organized as follows. In Section II, we will briefly review the line element and model the motion of massive and massless particles respectively. In Section III we model thin accretion disk in the geometry of the QFMGBH and use GYOTO, a general relativistic ray-tracing code Hua:2024jzx to simulate the image of the QFMGBH’s shadow. Finally in section IV we summarize the results and their implications. Throughout the article we use geometrized units setting $c=1$ with the metric signature ,+,+,+-,+,+,+.
II The spacetime and the motion of particles
@@
II.1 Black hole in quantum fluctuation modified gravity
In Hua:2024jzx , an interesting QFMGBH solution for the gravitational field equations in QFMG proposed in Yang:2015jla was obtained, which reduces to different classes of BHs surrounded by Kiselev fluids, by taking some specific values of the parameter of the equation of state (EoS). The geometry of the static and spherically symmetric QFMGBH in Schwarzschild coordinates is given by
$\begin{split}ds^{2}&=-g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}\\ &=-Bdt^{2}+\frac{1}{B}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2},\end{split}$ (1)
@@ II.1 Black hole in quantum fluctuation modified gravity
[…]
where $D$ is an integration constant, the parameter $\alpha$ characterizing the fluctuation of metric, and the parameter $\omega$ denotes the EoS of Kiselev fluid. If there are no fluctuations and fluids, solution (2) will reduce to the Schwarzschild BH. Here, we mainly focus on the impacts of parameters $\omega$ and $\alpha$ on the research objects. For the convenience of calculations, we take $D=\lambda M$. The dimension of $\lambda$ depends on the parameters $\omega$ and $\alpha$, so dose the physical property of the parameter $D$, see for example, when $\omega=0$ and $\alpha=0$, $D$ is equivalent to mass. When the BH is surrounded by different fluids, such as radiation, dust, quintessence, cosmological constant, or phantom, solution (2) reduces to different BH spacetime.
II.2 Asymptotically flat condition
@@
In the subsequent content, we will consider the Novikov-Thorne model for accretion, which bases on asymptotically flat spacetime. So we assume that the QFMGBH spacetime should be asymptotically flat at infinity, namely, it has to satisfy the condition $B\to 1$ when $r\to\infty$. In Yang:2015jla , the range of the parameter $\alpha$ is (1,1)(-1,1). We can further constrain $\alpha$ by the asymptotically flat condition, The relevant inequality is shown as
$P=-\frac{2(1+3\omega-4\omega\alpha)}{2-3\alpha+\omega\alpha}<0.$ (3)
Through analyzing the inequality (3), we can know that as $\alpha$ varies in range (1,1)(-1,1), the $P$ identically equals to $1$ for $\omega=-1$, which doesn’t satisfy the asymptotically flat condition. So we don’t consider this case ($\omega=-1$) further. In the same way, we can acquire the ranges of $\alpha$ for other different values of $\omega$: the range of $\alpha$ for $\omega=1/3$ is (1,3/4)(-1,3/4), for $\omega=0$ is (1,2/3)(-1,2/3), for $\omega=-2/3$ is $(3/8,6/11)$, and for $\omega=-4/3$ is $(6/13,9/16)$. We visualize these ranges in figure 1.
@@
II.3 Stable circular orbits
From the definition of the BH, we can know that the BH cannot be directly observed, but we can understand the properties of the central objects by researching the properties of its surroundings Urmanov_2024 , so the time-like geodesics around compact astrophysical objects are very crucial. As usually, we assume that the accretion disk is located in the equatorial plane of the BH ($\theta=\pi/2$). The Lagrangian of particles moving in the equatorial plane of a static spherically symmetric BH is as follows
$\begin{split}\mathcal{L}&=\frac{1}{2}g_{\mu\nu}\dot{x^{\mu}}\dot{x^{\nu}}\\ &=\frac{1}{2}g_{tt}\dot{t}^{2}+\frac{1}{2}g_{rr}\dot{r}^{2}+\frac{1}{2}g_{\phi\phi}\dot{\phi}^{2},\end{split}$ (4)
@@ II.3 Stable circular orbits
$\begin{split}g_{tt}\dot{t}&=-E,\\ g_{\phi\phi}\dot{\phi}&=L,\end{split}$ (5)
@@
$\begin{split}\dot{t}&=-\frac{E}{g_{tt}},\\ \dot{\phi}&=\frac{L}{g_{\phi\phi}}.\end{split}$ (6)
@@
$V_{\rm{eff}}(E,L,r)=-1+\frac{g_{\phi\phi}E^{2}+g_{tt}L^{2}}{-g_{tt}g_{\phi\phi}},$ (7)
where $V_{\rm{eff}}=g_{rr}\dot{r}^{2}$ denotes the effective potential of particles. Circular orbits exist at positions where $\dot{r}=0$, which corresponds to the conditions $V_{\rm{eff}}=0$ and $V_{\rm{eff}}^{\prime}=0$ Gair_2008 ; Bambi_2011(1) ; Bambi_2011(2) , where the prime denotes differentiation with respect to $r$. Under these conditions, we can derive expressions for the specific energy $E$, the specific angular momentum $L$, and the specific angular velocity $\Omega$ of particles in the equatorial plane of a Schwarzschild-like black hole spacetime.
$\Omega=\frac{d\phi}{dt}=\frac{\sqrt{-g_{tt,r}g_{\phi\phi,r}}}{g_{\phi\phi,r}},$ (8)
$E=-\frac{g_{tt}}{\sqrt{-g_{tt}-g_{\phi\phi}\Omega^{2}}},$ (9)
$L=\frac{g_{\phi\phi}\Omega}{\sqrt{-g_{tt}-g_{\phi\phi}\Omega^{2}}}.$ (10)
We can derive the innermost stable circular orbit (ISCO) for particles in the circular orbits by putting the Eqs. (8)–(10) into $V_{\rm{eff}}^{\prime\prime}=0$.
III Physical properties of thin accretion disks
@@
Firstly, we provide a brief overview of the key properties of the accretion disk in the Novikov-Thorne model Shakura:1972te ; Novikov:1973kta . $(\textbf{1})$ The spacetime around the central mass object is static, axisymmetric, and asymptotically flat. $(\textbf{2})$ The gravitational effect that the accretion disk generate can be ignored. $(\textbf{3})$ The accretion disk is geometrically thin and optically thick, with the disk’s semi-thickness $h$ being significantly smaller than its radial extent $r$. As a result, light is absorbed when it passes vertically through the disk. $(\textbf{4})$ The inner edge of the disk is defined by the radius of the marginally stable orbit, with particles confined between the ISCO and the outer edge. $(\textbf{5})$ The accretion disk coincides with the equatorial plane of the compact accreting object. $(\textbf{6})$ The radiation emitted from the disk is modeled as a black-body spectrum. $(\textbf{7})$ The mass accretion rate $\dot{M}$ is constant over time.
III.1 The radiant energy flux
Considering the accretion disk around a QFMGBH, the metric components as well as the accretion particles move in Keplerian orbits with a rotational velocity $\Omega$, the specific energy $E$, and the specific angular momentum $L$ depends only on the radial coordinate. The particles with the four-velocity $u^{\mu}$ form a disk which has an averaged surface density $\Sigma$
$\Sigma=\int^{h}_{-h}\langle\rho_{0}\rangle dz,$ (11)
where $h$ is the disk height during the time $\Delta t$, $\rho_{0}$ is the density of rest mass, and $\langle A\rangle$ is defined as the averaged value of $A$ in time $\Delta t$ with angle $2\pi$. The particles in the disk is modeled by an anisotropic fluid source which reads
$T^{\mu\nu}=\rho_{0}u^{\mu}u^{\nu}+2u^{(\mu}q^{\nu)}+t^{\mu\nu},$ (12)
where the stress tensor $t^{\mu\nu}$ and the energy flow vector $q^{\mu}$ are measured in the averaged rest-frame with $u_{\mu}q^{\mu}=0$. Here the specific heat is neglected. The torque of the disk structure is given by
$W^{r}_{\phi}=\int^{h}_{-h}\langle t^{r}_{\phi}\rangle dz.$ (13)
The most important physical measure recounting the disk is the radiation flux $F(r)$ over the disk surface, which can be obtained from the time and orbital average of the energy flux vector
$F(r)=\langle q^{z}\rangle.$ (14)
Using stress energy tensor $T^{\mu\nu}$ of the disk, we can define the four-vectors of the energy and of the angular momentum flux in a local basis, respectively, as
$E^{\mu}=-T^{\mu\nu}\left(\frac{\partial}{\partial t}\right)^{\nu},~~{\rm and}~~J^{\mu}=-T^{\mu\nu}\left(\frac{\partial}{\partial\phi}\right)^{\nu}.$ (15)
The structure equations of the disk could be obtained from the conservation laws of the rest mass, of the energy, and of the angular momentum. The conservation for the rest mass, $\nabla_{\mu}(\rho_{0}u^{mu})=0$, yields the time averaged rate of the rest mass accretion, which is independent of the disk radius
$\dot{M}_{0}\equiv-2\pi\sqrt{-g}\Sigma u^{r}={\rm constant},$ (16)
where dot denotes the differentiation with respect to the time coordinate. The conservation for the energy, $\nabla_{\mu}u^{\mu}=0$, yields
$\left[\dot{M}_{0}E-2\pi\sqrt{-g}\Omega W^{r}_{\phi}\right]_{,r}=4\pi\sqrt{-g}FE,$ (17)
where the comma denotes the derivative with respect to the radial coordinate. This is an equilibrium equation, implying that the energy transported by the rest mass flow, $\dot{M}_{0}E$, and the energy transported by the torque in the disk, $2\pi\sqrt{-g}\Omega W^{r}_{\phi}$, is well-adjusted by the energy emitted from the surface of the disk, $4\pi\sqrt{-g}FE$. The conservation for angular momentum, $\nabla_{\mu}J^{\mu}$, yields
$\left[\dot{M}_{0}L-2\pi\sqrt{-g}W^{r}_{\phi}\right]_{,r}=4\pi\sqrt{-g}FL,$ (18)
where the first term characterizes the angular momentum from the rest mass of the disk, the second term is the angular momentum transported by the torque in the disk, the right hand side term is the angular momentum conveyed from the disk’s surface by radiation. Utilizing the above two equations, we can obtain the expression for the energy flux $F(r)$ by eliminating $W^{r}_{\phi}$. Using the relation between the energy and the angular momentum, $dE=QdJ$, for which the circular geodesic orbits assumes the form $E_{,r}=\Omega L_{,r}$, we can find the flux $F(r)$ of the radiant energy over the disk in terms of the specific energy, of the angular momentum and of the angular velocity of the disk matter in spherical coordinates
$F(r)=-\frac{\dot{M_{0}}\Omega_{,r}}{4\pi\sqrt{-g/g_{\theta\theta}}(E-\Omega L)^{2}}\int_{r_{\rm{ISCO}}}^{r}(E-\Omega L)L_{,r}\,dr,$ (19)
where the tensor transformation from the original vertical flux $q^{z}$ (in cylindrical coordinates), $F(r)=\langle q^{z}\rangle=\langle q^{\theta}(r,\theta_{0})\sqrt{g_{\theta\theta}(r,\theta_{0})}\rangle$, when $z\ll r$ or $\theta_{0}\approx\pi/2$, has been used Collodel_2021 .